Here are some other application difficulties that I am encountering.
The equation D=1.2square root of h gives the distance, D, in miles that a person can see to the horizon from a height, h, in feet.
a. Solve this equation for h.
b. Long's Peak in the Rocky Mountain National Park, is 14,255 feet in elevation. How far can you see to the horizon from the top of Long's Peak? Can you see Cheyenne, Wyoming (about 89 miles away)?
PLEASE help.

Answer by nycsub_teacher(90)

We can write the equation like this:
D = 1.2(sqrt{h})
Divide both sides by 1.2 to isolate the radical.
D/1.2 = sqrt{h}
We now square both sides to remove the radical.
(D/1.2) = [sqrt{h}]^2
(D^2)/(1.44) = h
b.Long’s Peak in the Rocky Mountain National Park, is 14,255 feet in elevation. How far can you see to the horizon from the top of Long’s Peak? Can you see Cheyenne, Wyoming (about 89 miles away)?
For part b, let h = 14,255 and solve for D.

D = 1.2(sqrt{14,255})
D = 143.27 feet
Now convert this into miles. If it comes out to be greater than 89 miles, then the person can see Cheyenne, Wyoming, otherwise not.